# Conformalized Credal Regions for Classification with Ambiguous Ground Truth

### v1

@@ 5. Improving over Conformal Prediction Sets
This gives us an easy way to compute $\underline{P}(A)$ in practice, to then derive IDHSs. We can use (4) and Definition 1.(i) to derive Algorithm 1, a greedy algorithm to build $\text{IS}_{\mathcal{P},\delta}$. In turn, IHDS $\text{IS}_{\mathcal{P},\delta}$ is used to derive a predictive set narrower than the conformal one $\Psi(c(X_{n+1}))$ from (2), and that retains the same probabilistic guarantees. Recall that Definition 1.(i) gives us a probabilistic guarantee that holds for all distributions in the credal region $\mathcal{P}$ that we built in Corollary 4.1. This is a slightly weaker guarantee than that of classical conformal prediction, whose guarantee instead holds for all possible exchangeable distributions $P$ on $\mathcal{Y}$. Thanks to Proposition 5, though, a consequence of the following Proposition 7 is that the loss of coverage for the IHDS with respect to a classical CPS is negligible.
@@
Let $\alpha$ be the same significance level selected in (1). Pick any $\delta\in[0,1]$. Then, $\text{IS}_{\mathcal{P},\delta}\subseteq\Psi(c(X_{n+1}))$, and the inclusion is strict for some value of $\delta$. In addition, $\mathbb{P}[Y_{n+1}\in\text{IS}_{\mathcal{P},\delta}]\geq(1-\delta)(1-\alpha)$ where $Y_{n+1}$ denotes the correct label for input $X_{n+1}$.
@@
First, let us introduce the concept of $(1-\delta)$-(Precise) Highest Density Set $\text{HDS}_{P,\delta}$, for some $P\in\mathcal{P}$ [29],
@@
By [17, Page 3], we know that
@@
=$=\{y\in\mathcal{Y}:\exists P\in\mathcal{P}\text{, }P(y)\geq P^{\delta}\},$
@@
In standard conformal prediction – that is, when we calibrate on crisp-labeled data – the probabilistic guarantee that we derive is of the form $\mathbb{P}[Y_{n+1}\in\text{CPS}]\geq 1-\delta$, where we denote by CPS the conformal prediction set, and by $\delta$ the same threshold as in Proposition 7. In the present work, we need to take into account the imprecision coming from the ambiguous labeling, which appears in the form of $1-\alpha$ in the probabilistic guarantee of Proposition 7. In our imprecise probabilistic framework, $1-\alpha$ is the guarantee that we have on the true distribution $\textup{Cat}(\lambda_{n+1})$ belonging to the credal set $\mathcal{P}$ (see Proposition 5). But $1-\alpha$ is chosen by the user, so letting e.g. $0.951-\alpha=$ 0.95 or $0.991-\alpha=$ 0.99 will yield a negligible coverage loss with respect to the classical conformal prediction setting. It is also worth noting that the guarantee in Proposition 7 is the same derived for “standard” conformal prediction in the case of ambiguous labels in [53, Equation (43)].
@@
Step 1 For all $A\subseteq\mathcal{Y}$, compute $\underline{P}(A)$ using (4). $\triangleright$ $2^{K}$ possible $A$’s for $|\mathcal{Y}|=K$ labels.
@@
Finally, let us point out that we can use credal regions to quantify and disentangle between aleatoric and epistemic uncertainties (AU and EU, respectively) in the analysis at hand [8]. AU refers to the uncertainty that is inherent to the data generating process; as such, it is irreducible. EU, instead, refers to the lack of knowledge about the data generating process; as such, it is reducible. It can be lessened on the basis of additional data, e.g. by retraining the model using an augmented training set [36]. On the other hand, since AU is irreducible, there is an increasing need for ML techniques that are able to detect and flag excess of AU, so that the user can “proceed with caution”.
@@
The credal versions of the Shannon entropy as proposed by [1, 30] are $\overline{H}(P)\coloneqq\sup_{P\in\mathcal{P}}H(P)$ and $\underline{H}(P)\coloneqq\inf_{P\in\mathcal{P}}H(P)$, called the upper and lower (Shannon) entropy, respectively. The upper entropy is a measure of total uncertainty (TU), since it represents the minimum level of predictability associated with the elements of $\mathcal{P}$. In [1, 30], the authors postulate that TU can be decomposed additively as a sum of aleatoric and epistemic uncertainties, and that the latter can be specified as the difference between upper and lower entropy, thus obtaining
$\underbrace{\overline{H}(P)}_{\text{total uncertainty }\text{TU}(\mathcal{P})}=\underbrace{\underline{H}(P)}_{\text{aleatoric uncertainty }\text{AU}(\mathcal{P})}+\underbrace{\left[\overline{H}(P)-\underline{H}(P)\right].}_{\text{epistemic uncertainty }\text{EU}(\mathcal{P})}$ (6)
Other measures based on credal regions are also available (see [4, 27, 30] or Appendix A for a few examples) and they can be used in place of upper and lower entropy to quantify EU and AU within our credal region $\mathcal{P}$, as long as the measure chosen for the total uncertainty is bounded. We also note in passing that the decomposition in (6) is extremely important for the field of conformal prediction, since, as pointed out in [30, Section 5], the role of aleatoric and epistemic uncertainties in (classical) conformal prediction is in general not immediately clear. Our approach allows to overcome this shortcoming.
@@
6. Experiments
@@
We verify the proposed algorithm on three datasets with ambiguous ground truth, including the toy and Dermatology DDx datasets from [52, 53] (“Toy” and “Derm”, respectively), and CIFAR-10H [44] (“Cifar10h”). For Derm, we use risk labels as classes, classifying cases into low, medium and high risk. For CIFAR-10H, for computational convenience, we only consider data points whose annotated ground truth label in the original CIFAR-10 dataset is within the three classes (airplane, automobile, and bird). The toy dataset contains 1,000 data points, Dermatology DDx contains 1,947 data points, and CIFAR-10H contains 3,000 data points. We split each dataset into random calibration and testing sets in a 50%-50% ratio.
@@
We consider different miscoverage levels $\epsilon$, including $0.05$, $0.1$, $0.15$, $0.2$, $0.25$ and $0.3$. We denote the coverage level as $1-\epsilon$ in our plots. For simplicity, we put the significance levels $\alpha$ and $\delta$ to $\frac{\epsilon}{2}$. We compare our method (denoted by “Ours (IS)” for “Imprecise Set”) to Plausibility Reduced Predictive Set $\Psi(c(X_{n+1}))$ proposed in [53, Equation (41)] (denoted by “[49](PRPS)”). For Algorithm 1, we construct the credal region $c(X_{n+1})$ from (1) by discretizing the simplex and computing the convex hull. We report three measures, including empirical distribution coverage, empirical label coverage, and average inefficiency. To account for randomness, we run each experiment with 20 random seeds.
@@
Secondly, we report the end-to-end empirical label coverage levels in Figure 3. As shown in the figure, the true label coverage levels of both methods are above $1-\epsilon$, which is consistent with the coverage guarantee in [53, Equation (43)] and Proposition 7. Both our method and the baseline served to obtain the coverage guarantee.
Figure 2. Empirical distribution coverage levels. The left plot is on toy dataset; the middle plot is on Dermatology DDx dataset; the right plot is on CIFAR-10H.
@@
Figure 3. True label coverage levels. The left plot is on the toy dataset; the middle plot is on Dermatology DDx; the third is on CIFAR-10H.
Figure 4. Average inefficiencies. The left plot is on the toy dataset; the middle plot is on Dermatology DDx; the third is on CIFAR-10H.
@@
According to Proposition 7, for a specified coverage level $1-\epsilon$, the probability $\mathbb{P}[Y_{n+1}\in\text{IS}_{\mathcal{P},\delta}]$ holds for any combination of $\alpha$ and $\delta$ as long as they satisfy $(1-\delta)(1-\alpha)\geq 1-\epsilon$. However, different combinations of $\alpha$ and $\delta$ can impact the inefficiency of the resulting prediction set. To study this effect, we consider $\epsilon\in\{0.05,0.1,0.15,0.2,0.25,0.3,0.35,0.4\}$, select $\alpha$ on a grid in $(0,\epsilon)$ with intervals of $\epsilon/10$, and compute the corresponding $\delta$ levels. Figure 5 shows the average inefficiency across various $\alpha$ and $\delta$ combinations. The plot indicates that $\alpha$ has a more significant effect on inefficiency; allowing higher $\alpha$ typically results in lower inefficiency (i.e., smaller prediction sets).
Figure 5. Inefficiency using different combinations of $\alpha$ and $\delta$.
@@
Table 1. Qualitative Examples from the DDx Dataset: The left column presents the original medical images, the middle column shows the label distributions annotated by experts, and the right column displays the credal regions constructed using our method.
7. Conclusion
@@
In this paper, we address an important problem in imprecise probabilistic machine learning, namely how to empirically derive credal regions in a data-driven and efficient way, without any prior assumptions. To this end, we build on previous work by [52, 53] and apply conformal prediction in the probability space. Given a calibration set of examples with associated categorical distributions over classes, we can explicitly construct credal regions that enjoy a calibration property guaranteeing the true data generating process being included with high probability. Besides being efficient, this approach also allows to deal with ambiguous ground truth, examples where the ground truth label is uncertain. By constructing imprecise highest density sets [17], we can derive predictive sets of labels that are narrower, i.e., more efficient than those from [53]. Compared to the seminal work of [14, 15], we obtain credal sets without explicitly constructing the conformal transducer (i.e., computing $p$-values for all classes of the test example) and avoid the consonance assumption (stating that the $p$-value for the class $\tilde{y}$ that makes $(x_{n+1},\tilde{y})$ the “most conformal” to what we see in the calibration set has to be $1$). Moreover, due to our calibration property, we do not need to assume that the true data generating process is included in the credal region. We also show that our credal region is $[\delta/(1-\alpha)]$-type-2 valid, a version of a notion introduced in [15, Definition 2].
@@
Limitations. Eliciting $\mathcal{P}$ requires exchangeability as a consequence of using a conformal approach; it also requires the availability of a calibration set. In addition, some subjectivity enters the analysis via the choice of the non-conformity measure for the conformal prediction methodology. Furthermore, as proven e.g. in [33], with finite samples, we cannot give a conditional version of the probabilistic guarantee in Proposition 7. Finally, throughout the paper we (tacitly) assumed that the plausibility vectors $\lambda_{i}$ in the calibration set are available and accurate; this may not be the case in some applications [51].
Appendix A Uncertainty Quantification

### v2

@@ 4 Improving over Conformal Prediction Sets
This gives us an easy way to compute $\underline{P}(A)$ in practice, to then derive IDHSs. We can use equation 5 and Definition 1.(i) to derive Algorithm 1, a greedy algorithm to build $\text{IS}_{\mathcal{P},\delta}$. In turn, IHDS $\text{IS}_{\mathcal{P},\delta}$ is used to derive a predictive set narrower than the conformal one $\Psi(c(X_{n+1}))$ from equation 2, and that retains the same probabilistic guarantees. Recall that Definition 1.(i) gives us a probabilistic guarantee that holds for all distributions in the credal region $\mathcal{P}$ that we built in Corollary 4.1. This is a slightly weaker guarantee than that of classical conformal prediction, whose guarantee instead holds for all possible exchangeable distributions $P$ on $\mathcal{Y}$. Thanks to Proposition 5, though, a consequence of the following Proposition 7 is that the loss of coverage for the IHDS with respect to a classical CPS is negligible.
@@
Let $\alpha$ be the same significance level selected in equation 1. Pick any $\delta\in[0,1]$. Then, $\text{IS}_{\mathcal{P},\delta}\subseteq\Psi(c(X_{n+1}))$, and the inclusion is strict for some value of $\delta$. In addition, $\mathbb{P}[Y_{n+1}\in\text{IS}_{\mathcal{P},\delta}]\geq(1-\delta)(1-\alpha)$ where $Y_{n+1}$ denotes the correct label for input $X_{n+1}$.
@@
First, let us introduce the concept of $(1-\delta)$-(Precise) Highest Density Set $\text{HDS}_{P,\delta}$, for some $P\in\mathcal{P}$ Hyndman (1996),
@@
By (Coolen, 1992, Page 3), we know that
@@
=$=\{y\in\mathcal{Y}:\exists P\in\mathcal{P}\text{, }P(\{y\})\geq P^{\delta}\},$
@@
In standard conformal prediction – that is, when we calibrate on crisp-labeled data – the probabilistic guarantee that we derive is of the form $\mathbb{P}[Y_{n+1}\in\text{CPS}]\geq 1-\delta$, where we denote by CPS the conformal prediction set, and by $\delta$ the same threshold as in Proposition 7. In the present work, we need to take into account the imprecision coming from the ambiguous labeling, which appears in the form of $1-\alpha$ in the probabilistic guarantee of Proposition 7. In our imprecise probabilistic framework, $1-\alpha$ is the guarantee that we have on the true distribution $\textup{Cat}(\lambda_{n+1})$ belonging to the credal region $\mathcal{P}$ (see Proposition 5). But $1-\alpha$ is chosen by the user, so letting e.g. $0.951-\alpha=$ 0.95 or $0.991-\alpha=$ 0.99 will yield a negligible coverage loss with respect to the classical conformal prediction setting. It is also worth noting that the guarantee in Proposition 7 is the same derived for “standard” conformal prediction in the case of ambiguous labels in (Stutz et al., 2023c, Equation (43)).
@@
Step 1 For all $A\subseteq\mathcal{Y}$, compute $\underline{P}(A)$ using equation 5. $\triangleright$ $2^{K}$ possible $A$’s for $|\mathcal{Y}|=K$ labels.
@@
Finally, let us point out that we can use credal regions to quantify and disentangle between aleatoric and epistemic uncertainties (AU and EU, respectively) in the analysis at hand (Caprio et al., 2024b). AU refers to the uncertainty that is inherent to the data generating process; as such, it is irreducible. EU, instead, refers to the lack of knowledge about the data generating process; as such, it is reducible. It can be lessened on the basis of additional data, e.g. by retraining the model using an augmented training set (Lin et al., 2024). On the other hand, since AU is irreducible, there is an increasing need for ML techniques that are able to detect and flag excess of AU, so that the user can “proceed with caution”.
@@
The credal versions of the Shannon entropy as proposed by Abellán et al. (2006); Hüllermeier & Waegeman (2021) are $\overline{H}(\mathcal{P})\coloneqq\sup_{P\in\mathcal{P}}H(P)$ and $\underline{H}(\mathcal{P})\coloneqq\inf_{P\in\mathcal{P}}H(P)$, called the upper and lower (Shannon) entropy, respectively. The upper entropy is a measure of total uncertainty (TU), since it represents the minimum level of predictability associated with the elements of $\mathcal{P}$. In Abellán et al. (2006); Hüllermeier & Waegeman (2021), the authors postulate that TU can be decomposed additively as a sum of aleatoric and epistemic uncertainties, and that the latter can be specified as the difference between upper and lower entropy, thus obtaining
$\underbrace{\overline{H}(\mathcal{P})}_{\text{total uncertainty }\text{TU}(\mathcal{P})}=\underbrace{\underline{H}(\mathcal{P})}_{\text{aleatoric uncertainty }\text{AU}(\mathcal{P})}+\underbrace{\left[\overline{H}(\mathcal{P})-\underline{H}(\mathcal{P})\right].}_{\text{epistemic uncertainty }\text{EU}(\mathcal{P})}$ (7)
Other measures based on credal regions are also available (see Bronevich & Rozenberg (2021); Hofman et al. (2024); Hüllermeier & Waegeman (2021) or Appendix B for a few examples) and they can be used in place of upper and lower entropy to quantify EU and AU within our credal region $\mathcal{P}$, as long as the measure chosen for the total uncertainty is bounded. We also note in passing that the decomposition in equation 7 is extremely important for the field of conformal prediction, since, as pointed out in (Hüllermeier & Waegeman, 2021, Section 5), the role of aleatoric and epistemic uncertainties in (classical) conformal prediction is in general not immediately clear. Our approach allows to overcome this shortcoming.
That being said, there is an active ongoing debate around whether TU actually decomposes additively into AU and EU (Baan et al., 2023; Gruber et al., 2023; Kirchhof, 2024; Ulmer, 2024; Wimmer et al., 2023). Furthermore, Mucsányi et al. (2024) convincingly proved that uncertainty measures are very context- and task-specific, and so – even under the assumption of an additive decomposition – the choice of upper and lower entropy in equation 7 is subject to the specific analysis that the user is carrying out. Because of this, we defer to future work looking for the “correct” measures for AU and EU, and how such a choice depends on the problem at hand.
@@
It is worth noting that our method is not the only one exhibiting intrinsic UQ capabilities. Another notable class of such models are evidential deep learning ones (see e.g. Gao et al. (2024) for a comprehensive survey), in which uncertainty is modeled via second-order distributions, that is, using distributions over distributions. Second-order methods, though, have been recently shown to suffer from major pitfalls when used to quantify predictive EU due to their sensitivity to regularization parameters, and to underestimate predictive AU (Bengs et al., 2022; Pandey & Yu, 2023; Jürgens et al., 2024). In addition, they are not immediately comparable to credal methods because the imprecise probabilistic apparatuses that underpin them are inherently different. Indeed, contrary to second-order methods, credal regions are (convex and closed) collections of first-order probabilities.
Let us also add another comment. It is true that the results in Cella & Martin (2022a; b) could in principle be related to second-order methods. This because – through the consonance assumption $\max_{y\in\mathcal{Y}}\pi(y)=1$ – they effectively make the upper probability $\overline{\Pi}$ of the credal region $\mathcal{M}(\overline{\Pi})$ associated to conformal prediction a consonant plausibility function. In turn, the latter corresponds to a possibility function (Zadeh, 1986), which is a second-order distribution. In this paper, instead, we do not need consonance to derive our credal region, and so the connection with second-order approaches, if it exists, is more subtle. We defer looking for it to future work.
5 Experiments
@@
We verify the proposed algorithm on three datasets with ambiguous ground truth, including the toy and Dermatology DDx datasets from Stutz et al. (2023b; c) (“Toy” and “Derm”, respectively), and CIFAR-10H (Peterson et al., 2019b) (“Cifar10h”). For Derm, whose details are discussed in Appendix C, we use risk labels as classes, classifying cases into low, medium and high risk. For CIFAR-10H, for computational convenience, we only consider data points whose annotated ground truth label in the original CIFAR-10 dataset is within the three classes (airplane, automobile, and bird). The toy dataset contains 1,000 data points, Dermatology DDx contains 1,947 data points, and CIFAR-10H contains 3,000 data points. We split each dataset into random calibration and testing sets in a 50%-50% ratio. Furthermore, we use estimated classifier confidence scores as the conformity functions. Specifically, for the Toy dataset, we train a single-layer MLP with 100 hidden neurons, achieving an accuracy of 77.5%. For CIFAR-10H, we employ a ResNet50 (He et al., 2015) model trained on the original CIFAR-10 training set, obtaining an accuracy of 93.6%. For the Derm dataset, we rely on the confidence scores collected from the testing set reported by (Stutz et al., 2023c).
[…]
@@
We consider different miscoverage levels $\epsilon$, including $0.05$, $0.1$, $0.15$, $0.2$, $0.25$ and $0.3$. We denote the coverage level as $1-\epsilon$ in our plots. For simplicity, we put the significance levels $\alpha$ and $\delta$ to $\frac{\epsilon}{2}$. We compare our method (denoted by “Ours (IS)” for “Imprecise Set”) to Plausibility Reduced Predictive Set $\Psi(c(X_{n+1}))$ proposed in (Stutz et al., 2023c, Equation (41)) (denoted by “PRPS”). For Algorithm 1, we construct the credal region $c(X_{n+1})$ from equation 1 by discretizing the simplex and computing the convex hull. We report three measures, including empirical distribution coverage, empirical label coverage, and average inefficiency. In particular, let $\{\lambda_{1},\ldots,\lambda_{N}\}$ be the annotated label distributions on the $N$ testing datapoints, $\{\mathcal{P}_{1},\ldots,\mathcal{P}_{N}\}$ be the constructed credal regions, $\{\text{IHDS}_{1},\ldots,\text{IHDS}_{N}\}$ be the constructed Imprecise Highest Density Sets. The empirical distribution coverage is defined as
$\frac{1}{N}\sum_{n=1}^{N}\mathbb{I}[\lambda_{n}\in\mathcal{P}_{N}],$ (8)
where $\mathbb{I}[\cdot]$ denotes the indicator function. The empirical label coverage is defined as
$\frac{1}{N}\sum_{n=1}^{N}\sum_{k=1}^{K}\mathbb{I}[k\in\text{IHDS}_{N}]\times\lambda_{N}^{k},$ (9)
where $\lambda_{N}^{k}$ is the annotated probability for $k$-th class. The average inefficiency is defined as
$\frac{1}{N}\sum_{n=1}^{N}|\text{IHDS}_{n}|,$ (10)
where $|\text{IHDS}_{n}|$ is the cardinality of the $n$-th IHDS. To account for randomness, we run each experiment with $20$ random seeds.
@@
Secondly, we report the end-to-end empirical label coverage levels in Figure 3. As shown in the figure, the true label coverage levels of both methods are above $1-\epsilon$, which is consistent with the coverage guarantee in (Stutz et al., 2023c, Equation (43)) and Proposition 7. Both our method and the baseline served to obtain the coverage guarantee. Furthermore, the empirical coverage is higher than the expected level because, even if the true distribution is not fully contained within the specified credal region, the resulting high density set might still include the true label. This mechanism elevates the empirical coverage above $(1-\alpha)(1-\delta)$.
Figure 2: Empirical distribution coverage levels. The left plot is on toy dataset; the middle plot is on Dermatology DDx dataset; the right plot is on CIFAR-10H.
@@
Let us add a remark on the differences between the datasets. They stem from the underlying ambiguity of the datasets. Compared to CIFAR-10H, the Dermatology DDX dataset is significantly more challenging by having more ambiguity in the ground truth labels. This produces sets that are more conservative overall, leading to higher coverage at lower variance.
Figure 3: True label coverage levels. The left plot is on the toy dataset; the middle plot is on Dermatology DDx; the third is on CIFAR-10H.
Figure 4: Average inefficiencies. The left plot is on the toy dataset; the middle plot is on Dermatology DDx; the third is on CIFAR-10H.
@@
According to Proposition 7, for a specified coverage level $1-\epsilon$, the probability $\mathbb{P}[Y_{n+1}\in\text{IS}_{\mathcal{P},\delta}]$ holds for any combination of $\alpha$ and $\delta$ as long as they satisfy $(1-\delta)(1-\alpha)\geq 1-\epsilon$. However, different combinations of $\alpha$ and $\delta$ can impact the inefficiency of the resulting prediction set. To study this effect, we consider $\epsilon\in\{0.05,0.1,0.15,0.2,0.25,0.3,0.35,0.4\}$, select $\alpha$ on a grid in $(0,\epsilon)$ with intervals of $\epsilon/10$, and compute the corresponding $\delta$ levels. Figure 5 shows the average inefficiency across various $\alpha$ and $\delta$ combinations. The plot indicates that $\alpha$ has a more significant effect on inefficiency; allowing higher $\alpha$ typically results in lower inefficiency (i.e., smaller prediction sets). Furthermore, given a $(1-\alpha)(1-\delta)$ confidence level, variations in $\alpha$ have a more pronounced impact on the resulting inefficiency. Therefore, we recommend allocating more error tolerance to $\alpha$ while maintaining a reasonable $\delta$. Developing a technique to identify the optimal combination of $\alpha$ and $\delta$ levels remains a direction for future work.
Figure 5: Inefficiency using different combinations of $\alpha$ and $\delta$.
@@
In addition, we also compute lower and upper entropy for both examples. For the first one, we have that $\text{AU}(\mathcal{P})=\underline{H}(\mathcal{P})=0$ and $\text{TU}(\mathcal{P})=\overline{H}(\mathcal{P})=1.091$, so that $\text{EU}(\mathcal{P})=\overline{H}(\mathcal{P})-\underline{H}(\mathcal{P})=1.091$. For the second one, we have that $\text{AU}(\mathcal{P})=\underline{H}(\mathcal{P})=0$ and $\text{TU}(\mathcal{P})=\overline{H}(\mathcal{P})=0.991$, so that $\text{EU}(\mathcal{P})=\overline{H}(\mathcal{P})-\underline{H}(\mathcal{P})=0.991$. As we can see, in both cases the entire uncertainty faced by the user is of epistemic nature. The reason why this happens is that in both cases at least one one-hot encoding probability vector belongs to the credal region $\mathcal{P}$. In particular, in both cases the credal regions contain the one-hot encoding vector for the label “Medium”, $\lambda^{\star}=(0,1,0)^{\top}$. Then, given the current state of epistemic knowledge, we cannot exclude that if we were to collect enough extra data, we would end up finding a singleton credal region $\mathcal{P}=\{P^{\star}=\text{Cat}(\lambda^{\star})\}$. In that case, we would have zero aleatoric uncertainty, since we would be sure (with $P^{\star}$-probability $1$) that the label “Medium” is the correct one for the new input. In turn, since for the time being we cannot exclude this best-case scenario, all the uncertainty encoded in the credal regions is of reducible nature (EU).
@@
Table 1: Qualitative Examples from the DDx Dataset: The left column presents the original medical images, the middle column shows the label distributions annotated by experts, and the right column displays the credal regions constructed using our method.
6 Conclusion
@@
In this paper, we address an important problem in imprecise probabilistic machine learning, namely how to empirically derive credal regions in a data-driven and efficient way, without any prior assumptions. To this end, we build on previous work by Stutz et al. (2023b; c) and apply conformal prediction in the probability space. Given a calibration set of examples with associated categorical distributions over classes, we can explicitly construct credal regions that enjoy a calibration property guaranteeing the true data generating process being included with high probability. Besides being efficient, this approach also allows to deal with ambiguous ground truth, examples where the ground truth label is uncertain. By constructing imprecise highest density sets (Coolen, 1992), we can derive predictive sets of labels that are narrower, i.e., more efficient than those from Stutz et al. (2023c). Compared to the seminal work of Cella & Martin (2022a; b), we obtain credal regions without explicitly constructing the conformal transducer (i.e., computing $p$-values for all classes of the test example) and avoid the consonance assumption (stating that the $p$-value for the class $\tilde{y}$ that makes $(x_{n+1},\tilde{y})$ the “most conformal” to what we see in the calibration set has to be $1$). Moreover, due to our calibration property, we do not need to assume that the true data generating process is included in the credal region. We also show that our credal region is $[\delta/(1-\alpha)]$-type-2 valid, a version of a notion introduced in Cella & Martin (2022b, Definition 2).
@@
Limitations. Eliciting $\mathcal{P}$ requires exchangeability as a consequence of using a conformal approach; it also requires the availability of a calibration set. In addition, some subjectivity enters the analysis via the choice of the non-conformity measure for the conformal prediction methodology. Moreover, as proven e.g. in Lei & Wasserman (2013), with finite samples, we cannot give a conditional version of the probabilistic guarantee in Proposition 7. Furthermore, the discretization involved in constructing credal regions can lead to high computational complexity, particularly when the label space is large. Finally, throughout the paper we (tacitly) assumed that the plausibility vectors $\lambda_{i}$ in the calibration set are available and accurate; this may not be the case in some applications (Stutz et al., 2023a).
Appendix A Differences Between $\mathbb{P}_{\text{agg}}$ and $\mathbb{P}$
Following the argumentation of Stutz et al. (2023b, Section 3.1), the first step is to realize that in reality, we always deal with $\mathbb{P}_{\text{agg}}$ rather than $\mathbb{P}$. This is because in most cases, our labels are derived from annotations rather than from some objective process. Even if a ground truth is obtained from a measurement not involving human annotation, there might be noise and errors in these measurements. With this realization, the best we can do in terms of a coverage guarantee is to guarantee coverage with respect to $\mathbb{P}_{\text{agg}}$ (since we do not know and will likely never know the true underlying $\mathbb{P}$). In other words, performing well against $\mathbb{P}_{\text{agg}}$ means performing well against expert annotators. In light of $\mathbb{P}$ being unavailable, performing as good as the expert annotations is the best we can ever hope for.
Mathematically, we can model $\mathbb{P}_{\text{agg}}$ using an annotation process as described in detail in the last two paragraphs of Stutz et al. (2023b, Section 3.1).
Appendix B Uncertainty Quantification